a shot is fired from a point at a distance of 200m from the foot of tower 100m high. so that it just passes over it. find the direction of shot

Grade:11

3 Answers

Vikas TU
6 years ago
Dear Student,

It just passes over the wall means the maximum height of the projectile is equal to the height of the wall.
hmax=100
Range=2x Distance between point of projection and the tower
=400m
By putting the formula of range and height we know that,
hmax/Range=tan
q/4
=>tan
q=1
=>
q=45⁰

The shot is fired at an agle of 45⁰ with the horizontal in the direction of the wall.

Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
gavini haritha
6 years ago
Given r=200m,and maximum height is 100m. Draw a triangle from point of projection to h, r/2. Then r/2=100 , tanteta=100/100=1 i.e 45
hemanth
6 years ago
r=200
hmax=100
r/2=100
triangle with opp side h
adj side r/2 point iof projection
thus tan teeta =100/100=1
thus tan teeeta =1
thus teeta =45 deg